# A cosmological “Big Storm Scenario" following the QCD phase transition

### v1

@@ II Soft graviton emission without a storm
where $n_{a},n_{b}$ are densities of colliding particles, $V$ is interacting volume, $v$ velocity and $\sigma_{transport}$ is the transport cross section (weighted with $sin(\theta)^{2}$). As an exercise, for electron-proton collisions in the Sun, he estimated this power. Substituting regulated cross section one gets the power of the Sun about $P_{Sun\,gravitons}\approx 70MW$ emitted in sub-$keV$ gravitons.
For comparison, total Sun power in thermal photons is $P_{Sun}=4.4\cdot 10^{16}Watts$. A reader may wonder why GW radiation is only billion times smaller than electromagnetic one, in spite of the fact that its coupling is smaller by about 38 orders of magnitude? Without going into complicated nuclear physics of the Sun energy production, let us just note one effect: soft photons are emitted into a heat bath in thermal equilibrium, so their emission must be balanced by photon absorption. GWs, on the other hand, are never close to thermal equilibrium, inverse reactions are negligible and their radiation proceeds unobstructed as estimated.
Following this reasoning, let us evaluate the production of soft gravitons accompanying hadronic collisions in hadron resonance gas before/after the QCD phase transition. In the plasma phase, effective masses of quark and gluon quasiparticles are $O(1\,GeV)$. Most hadronic species in HRG have the same order of magnitude. Comparing to typical thermal momenta $3T_{QCD}\sim 0.5\,GeV$ one may take the nonrelativistic estimate as qualitatively valid. With $M=1\,GeV,n_{a}=n_{b}=1/fm^{3},V=1\cdot m^{3}$ we estimate soft graviton power production to be
@@
The energy produced during time $t_{QCD}\sim 10^{-5}s$ gives the energy density (per $m^{3}$)
$\rho_{gravitons}\sim 1.6\cdot 10^{24}{kg\over ms^{2}}$
For comparison, the energy density of matter itself is
$\rho_{matter}\sim 1\,GeV/fm^{3}\sim 1.6\cdot 10^{35}{kg\over ms^{2}}$
eleven orders of magnitude larger. A very small radiation rate, containing a very small dimensionless ratio (4), is compensated by the relatively long duration of the emission time (the QCD era): gravitons, unlike photons and strongly interacting species, are only emitted, with negligible absorption. This type of enhancement of any “penetrating probe" production is well known in the heavy-ion collision community, among others.
A more formal answer to the question of why it is not suppressed much more can be given as follows: since the duration $t_{QCD}$ (1) is proportional to the Plank mass $M_{P}=\sqrt{\hbar\cdot c/G_{N}}$, the energy density of the soft GW produced is in fact proportional to a square root $\sqrt{G_{N}}$, not small coupling $G_{N}$ itself.
There are recent studies of soft graviton production from cosmological perturbations Sharma et al. (2023). These authors did find a universal “shallow relic" of the acoustic peak at small momenta, but that has an energy spectrum linear in momentum $\Omega_{GW}\sim k$. Their model is very different from soft radiation from particle collisions considered above, and thus, there is no contradiction between these results.
It is convenient to convert the observed radiation in energy density units at the current time to the time $t_{QCD}$. If this radiation is primordial, its energy density scales as $z^{4}$. The corresponding factor $z_{QCD}=T_{QCD}/T_{now}\approx 6\cdot 10^{11}$. Using observed
@@
we get
$\rho_{GW\,QCD}\approx 9\cdot 10^{30}{kg\over ms^{2}}$ (8)
This is of the order of $10^{-5}$ of matter density, but also six orders of magnitude larger than due to soft graviton emission evaluated above.
@@ III Acoustic turbulence as an ensemble of (relativistic) shocks
The turbulence of acoustic waves in a medium with sufficiently low viscosity is a set of shock waves, see e.g. Fig.1 taken from the recent paper Kochurin, E.A. and Kuznetsov,E.A (2022).
@@
In our case of QGP/hadronic gas, the velocity of even weak shocks coincides with the speed of sound $c_{s}=c/\sqrt{3}\approx 0.55\cdot c$, comparable to the speed of light. Stronger shocks move even with a larger speed. Therefore, one must use relativistic notations, so we explain those first. Hydrodynamics is described by motion of matter defined via (Lorentz scalars) energy density $\rho$ and pressure $p$, and the collective flow velocity. It is promoted to 4-vector $u^{\mu},\mu=0,1,2,3$ making the stress tensor in local approximation (zero viscosity) as
@@
Even if the motion happens in one dimension, there are two components of 4-velocity $u^{0},u^{1}$, constrained by $(u^{0})^{2}-(u^{1})^{2}=1$. Using $u^{0}=cosh(y),u^{1}=sinh(y)$ one uses a parameter $y$ called $rapidity$. Unlike velocity, its Lorentz transformation is done by simple additive shift.
@@
With increasing $Y$ the compression contrast between upstream and downstream matter grows, reaching about an order of magnitude for $Y\sim 1$. Even for more realistic $Y\approx 1/2$, the shock rapidity $y_{shock}\approx 1$ and velocity $v_{shock}=tanh(y)\approx 0.76$, and compression is already $\sim 2.5$.
The paper is Shuryak (2012) mostly devoted to calculation of the profile of strong shocks in strongly coupled plasma, using AdS/CFT correspondence. It may be used for more detailed studies of shock collisions and GW radiation, but not it is not used in the present paper and thus not discussed.
@@ IV Gravity wave emission from shock ensemble
We suggest that a “universal shoulder" in the spectrum toward the IR appears due to the emission of soft gravitons. As we discussed above, it is due to general expressions for soft emissions from collision processes in which the stress tensor suffers rapid changes. An important question is how far the soft GW shoulder reaches toward small $k$. Note that their emission happens during the time between collisions. As pointed out by Feinberg and Pomeranchuk in 1950’s, after each scattering, the electron appears “semi-naked", with a reduced electric field, which gradually regenerates later. Obviously, the same should be true for soft GW.
If shocks move relativistically, with Lorentz factor significantly larger than 1, their gravity field also gets Lorentz contracted into a pancake. When they change velocity or direction of motion, part of their Newtonian field continues into the old direction, constituting the soft radiation we discuss. We, therefore, state that the universal shoulder should reach the mean free path of the radiated objects. For the ensemble of shocks this longest time scale is
@@ V Primordial black holes from shock collisions?
Returning to the scenario under consideration, one may ask if colliding shocks may provide strong enough density perturbation for a gravitational collapse. While locally the shocks themselves include significant matter compression, their sizes still are expected to be much smaller than visible size of the Universe, $d,R\ll c\cdot t_{QCD}$. On the other hand, their collisions include matter going toward the center with relativistic rapidities, which undoubtedly helps to initiate the gravitational collapse. Corresponding calculations can be done by simulating Einstein equation in such geometry, which is a task for subsequent works.
Ultrarelativistic collisions of shock-like objects in the hydrodynamical framework were pioneered by Landau Landau (1953), who applied hydrodynamics to ultrarelativistic collisions of hadrons and nuclei. Due to Lorentz contraction, those look like “pancakes of matter". While considered crazy at the time, it does described well the first rapidity distributions from colliders in 1970’s Shuryak (1972). Now it is proven that relativistic hydrodynamics does provide quantitative description of heavy ion collisions, in many details see e.g. Shuryak (2017).
A revival of studies of ultrarelativistic collisions happened in the framework of AdS/CFT holographic correspondence, relating certain strongly coupled gauge theories to string theories in higher dimensions. In the so called Maldacena limit, the latter basically being the Einstein’s general relativity in 5 dimensions. Such correspondence reproduces many observed features of Quark-Gluon Plasma (QGP) including the equation of state and viscosity. Ultrarelativistic collisions of shock-like objects were studied numerically and analytically, see e.g. Chesler and Yaffe (2011) and subsequent literature.
We want to emphasize one aspect of such calculations clarifying the production of the BHs in shock collisions. When colliding objects are about to overlap for the first time, so that the nontrivial solution of Einstein equations has not yet started, one can already study if the so-called trapped surface (for massless particles) may or may not be formed. If it does, its area provides the lower bound for the BH entropy. This tool was suggested by Gubser and used in several publications a decade ago. Most interesting feature of the trapped surface (found numerically by one of us in Lin and Shuryak (2009) and demonstrated in Fig.4 from Gubser et al. (2009)) is sharp disappearence of the trapped surface. Shown in the figure is the trapped surface $area$ as a function of the impact parameter. At a critical value of the impact parameter, the trapped surface disappears, while the area is still nonzero. That implies that the mass distribution of produced BHs may have a qute sharp cutoff at some minimal value.

### v2

@@ II Soft graviton emission without a storm
where $n_{a},n_{b}$ are densities of colliding particles, $V$ is interacting volume, $v$ velocity and $\sigma_{transport}$ is the transport cross section (weighted with $sin(\theta)^{2}$). As an exercise, for electron-proton collisions in the Sun, he estimated this power. Substituting regulated cross section one gets the power of the Sun about $P_{Sun\,gravitons}\approx 7*10^{7}\,W$ emitted in sub-$keV$ gravitons. For comparison, total Sun power in thermal photons is $P_{Sun}=3.8\cdot 10^{26}\,W$, and comparable in solar neutrinos.
(A reader may wonder why GW radiation is only 19 orders of magnitude smaller than electromagnetic, while its coupling is smaller than $e^{2}/\hbar c=1/137$ by about 38 orders of magnitude? Without going into complicated nuclear physics of the Sun energy production, let us just note one effect: soft photons are emitted into a heat bath in thermal equilibrium, with emission balanced by absorption. GW radiation, on the other hand, just leave unobstructed.)
Let us now evaluate production of soft gravitons accompanying hadronic collisions in hadron resonance gas (HRG) after the QCD phase transition. In the plasma phase, effective masses of quark and gluon quasiparticles are $O(1\,GeV)$. Most hadronic species in HRG have the same order of magnitude. Comparing to typical thermal momenta $3T_{QCD}\sim 0.5\,GeV$ one may take the nonrelativistic estimate as qualitatively valid. With $M=1\,GeV,n_{a}=n_{b}=1/fm^{3},V=1\cdot m^{3}$ we estimate the soft graviton production to have power per $dVdt$
@@
The energy of GW produced during time $\sim 10t_{QCD}\sim 10^{-4}s$ is accumulated gives the achieved energy density (per $m^{3}$)
$\rho_{gravitons}\sim 6\cdot 10^{14}{kg\over ms^{2}}$
For comparison, order of magnitude of the energy density of hadronic matter itself is of the order of
$\rho_{hadrons}\sim 1\,GeV/fm^{3}\sim 1.6\cdot 10^{35}{kg\over ms^{2}}$
is 21 orders of magnitude larger.
(Note that small radiation rate, containing a very small dimensionless ratio (4), is somewhat compensated by the relatively long duration of the emission time. Gravitons, unlike photons and strongly interacting species, are accumulated during emission time, with negligible absorption. This type of enhancement of any “penetrating probe" production is well known in the heavy-ion collision community, among others. A more formal answer is to note that the duration $t_{QCD}$ (1) is proportional to the Plank mass $M_{P}=\sqrt{\hbar\cdot c/G_{N}}$, so the total energy density of the soft GW produced is in fact proportional to a square root $\sqrt{G_{N}}$, not the coupling $G_{N}$ itself.)
In recent studies of soft GW production from cosmological perturbations Sharma et al. (2023) a universal shoulder in spectrum left of the acoustic peak was observed, called “shallow relic" by the authors. Its energy spectrum is not constant but linear in momentum $\Omega_{GW}\sim k$: the difference with our shoulder is because their model is quite different, so there is no contradiction between both results.
After our simiplest estimates, let us jump to pulsar data and see how much GW power is needed to reproduce them. The observed radiation in energy density units now can be converted to the phase transition time $t_{QCD}$ by energy density scaling $z^{4}$. The corresponding factor $z_{QCD}=T_{QCD}/T_{now}\approx 6\cdot 10^{11}$. Using observed
@@
we get the energy density at the time of the QCD transition
$\rho_{GW\,then}\approx 9\cdot 10^{30}{kg\over ms^{2}}$ (8)
This is only $10^{-5}$ of matter density, but many orders of magnitude larger than due to soft graviton emission from hadron collisions evaluated above. Also of course the frequency do not match. So, the simplest source of GW production – hadron collisions – fails badly.
A “storm" idea substitute microscopic colliding objects (hadrons) by macroscopic ones, the “pancakes" of shocks, so masses and cross sections are greatly increased. The shoulder may reach correct wavelengths of $km$. At this time we cannot provide quantitative estimates of those, so we do not yet know if this mechanism can or cannot explain the observations.
@@ III Acoustic turbulence as an ensemble of (relativistic) shocks
The turbulence of acoustic waves in a medium with sufficiently low viscosity leads to a cascade of sound waves, eventually evolving into a set of shock waves, see e.g. Fig.1 taken from recent numerical simulation Kochurin, E.A. and Kuznetsov,E.A (2022).
@@
In our case of QGP/hadronic gas, the velocity of even weak shocks coincides with the speed of sound $c_{s}=c/\sqrt{3}\approx 0.55\cdot c$, and is comparable to the speed of light. Stronger shocks move even with a larger rapidities. Therefore, one must use relativistic notations we are going to remind now. Hydrodynamics is about motion of matter defined via energy density $\rho$, pressure $p$, and collective flow velocity vector $v_{i},i=1,2,3$. Two former quantities are Lorentz scalar, and velocity is promoted by 4-vector $u^{\mu},\mu=0,1,2,3$. The stress tensor in local approximation is
@@
Consider motion in one dimension, so there are two components of 4-velocity $u^{0},u^{1}$, constrained by $(u^{0})^{2}-(u^{1})^{2}=1$. Using $u^{0}=cosh(y),u^{1}=sinh(y)$ one introduced the parameter $y$ called $rapidity$. Unlike velocity, its Lorentz transformation is done by a simple shift.
@@
With increasing $Y$ the compression contrast between upstream and downstream matter grows, reaching about an order of magnitude for $Y\sim 1$. As a numerical example to be mentioned below, for $Y\approx 1/2$ the shock rapidity $y\approx 1$ and velocity $v_{shock}=tanh(y)\approx 0.76$, and compression $\sim 2.5$.
@@ IV Gravity wave emission from shock ensemble
We claim that a universal shoulder in the spectrum toward the IR is due to the emission of soft gravitons. As we discussed above, it is due to general expressions for soft emissions from collision processes in which the stress tensor suffers rapid changes.
An important question is how far the soft GW shoulder reaches toward small $k$. Note that their emission happens during the time between collisions. As pointed out by Feinberg and Pomeranchuk in 1950’s, after each scattering, the electron appears “semi-naked", with a reduced electric field, which gradually regenerates later. Obviously, the same should be true for soft GW.
If shocks move relativistically, their gravity field also gets Lorentz contracted into a pancake. When they change velocity or direction of motion, part of their Newtonian field continues into the old direction, constituting the soft radiation we discuss. We, therefore, state that the universal shoulder should reach the mean free path of the radiated objects. For the ensemble of shocks this longest time scale is
@@ V Primordial black holes from shock collisions?
Returning to the scenario under consideration, one may ask if colliding shocks may provide strong enough density perturbation for a gravitational collapse. While locally the shocks themselves include significant matter compression, their sizes still are expected to be much smaller than visible size of the Universe, $d,R\ll c\cdot t_{QCD}$. On the other hand, their collisions include matter going toward the center with relativistic rapidities, which undoubtedly helps to initiate the gravitational collapse. Corresponding calculations can be done by simulating Einstein equation in such geometry, which is a task for the future.
Ultrarelativistic collisions in the hydrodynamical framework were pioneered by Landau Landau (1953), who applied hydrodynamics to ultrarelativistic collisions of hadrons and nuclei. Due to Lorentz contraction, those look like “pancakes of matter". While considered crazy at the time of writing, it described well the first rapidity distributions from colliders Shuryak (1972). Now, the relativistic hydrodynamics does provide a description of heavy ion collisions in many details, see e.g. Shuryak (2017).
A revival of studies of ultrarelativistic collisions happened in the framework of AdS/CFT holographic correspondence, relating certain strongly coupled gauge theories to string theories in higher dimensions. In the so called Maldacena limit, the latter basically are Einstein’s general relativity in 5 dimensions. Such correspondence reproduces many observed features of Quark-Gluon Plasma (QGP) including the equation of state and viscosity. Ultrarelativistic collisions of shock-like objects were studied numerically and analytically, see e.g. Chesler and Yaffe (2011) and subsequent literature.
We want to emphasize one aspect of such calculations clarifying the production of the BHs in shock collisions. When colliding objects are about to overlap for the first time, so that the nontrivial solution of Einstein equations has not yet started, one can already study if the so-called trapped surface (for massless particles) may or may not be formed. If it does, its area provides the lower bound for the BH entropy. This tool was used in several publications a decade ago. A very interesting feature of the trapped surface (found numerically by one of us in Lin and Shuryak (2009)) is demonstrated in Fig.4 from Gubser et al. (2009). What is shown is the trapped surface area as a function of the impact parameter. At some critical value of the impact parameter, the trapped surface disappears completely while the area is still nonzero. That implies that mass distribution of produced BHs may have a sharp cutoff at some nonzero value.